An employee who produces x units per hour earns an hourly wage of y = 0.35x + 11 (in dollars). Find the hourly wage for an employee who produces 2 units per hour.



A) $11.40 B) $12.10 C) $12.20 D) $11.70


Answer by checkley71(8403)

Y=.35X+11 OR Y=.35*2+11 OR Y=.7+11 OR Y=11.7 OR $11.70 PER HOUR